OpenStudy (anonymous):

A physicist found that a force of 0.68 N was measured between two charged spheres. The distance between the spheres was 1.0 m. Calculate the force between the two spheres when the distance was decreased to 0.50 m. F2 = _____ N

OpenStudy (anonymous):

Using Coulomb's Law, F= ke*(q1*q2)/r^2 for the force between 2 charged spheres. where ke is a proportionality constant, q1 and q2 are the charges of sphere 1 and sphere 2 respectively and r is the distance in meters between the centers of both spheres. Give 0.68 N for a distance of 1.0m, 0.68 N = (ke*q1*q2)/ (1.0)^2 ke*q1*q2 = 0.68 Nm^2 [Using this as a constant since they're not changing]* For the new Force F2 = (ke*q1*q2)/(0.5^2) = 0.68 N m^2 /0.5^2 = 2.72N ke is a constant with a value of 8.988 N m^2 C^-2

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

No problemo ;D